OpenStudy (anonymous):

Write the equation of a circle with a center at (0, 3) and a radius of 1. x^2 + (y + 3)^2 = 1 <-- I came up with this am I correct? @waterineyes

OpenStudy (anonymous):

Slight mistake..

OpenStudy (anonymous):

is it x^2 + (y-3)^2 =1

OpenStudy (anonymous):

Equation of Circle is given by: \[(x - h)^2 + (y - k)^2 = r^2\]

OpenStudy (anonymous):

Yeah this one is just fine... Well Done..

OpenStudy (anonymous):

Thank you. Now I did another one. Would you mind telling me if I am correct on this one? Write the equation of a circle with a center at (–3, –8) and a radius of 1/3 I got (x + 3)^2 + (y + 8)^2 = 1/3

OpenStudy (anonymous):

See on right hand side it is radius square.. So earlier it was 1 so you did 1^2 = 1. But now it is 1/3, so what is the square of 1/3 ?? That you will write on RHS..

OpenStudy (anonymous):

Is it 1/9?

OpenStudy (anonymous):

In equation, on RHS, we write r^2 ie radius square, so you can't write 1/3 instead write square of it..

OpenStudy (anonymous):

Yeah that is right...

OpenStudy (anonymous):

Thank you! You have been an amazing help!

OpenStudy (anonymous):

No, I just know fewer things and not at all good at math... :)

OpenStudy (anonymous):

You are welcome dear.. Keep practicing more, Practice is other name of Math.. :P